OpenStudy (anonymous):

Recall that the sum of the measures of the interior angles of a triangle is 180 degrees. in the triangle of interest, two of the interior angles have the same measure, while the third interior angle measures 42 degrees less than each of the equivalent interior angles. find the measures of each interior angles.

OpenStudy (anonymous):

2x+x-42=180 3x-42=180 3x=222 x=222/3=74

OpenStudy (anonymous):

Let the angles be A, B and C. Let A = B. C = A - 42 C = B - 42 We all know A, B and C total must be 180. That should help figure out C and therefore A and B.

OpenStudy (anonymous):

answer is quite obvious 3x-42 =222 x = 74

OpenStudy (anonymous):

It's like what maxbar said, 74 is equal to the two equal angles. 32 is the smaller one (74-42).

OpenStudy (anonymous):

sorry gt bro gave the solution

OpenStudy (anonymous):

oo sardarji aisa mat karo

OpenStudy (anonymous):

or you could just do 180-42 and then divide taht by 2